OpenStudy (anonymous):

6x^3-18x^2-5x+15 factor by grouping

OpenStudy (anonymous):

3x-5x+15???

OpenStudy (anonymous):

https//webmath.com//

OpenStudy (anonymous):

(3x-5)(x+15) is this right?

OpenStudy (anonymous):

Well, the instruction tells you what to do. Note that the ratio between the first and second terms is the same as the ratio between the third and fourth. That is the indicator that grouping will work as a factoring strategy.\[6x^3-18x^2-5x+15=6x^2(x-3)-5(x-3)=(6x^2-5)(x-3)\]If we want to continue factoring, we can use difference of squares to get\[(6x^2-5)(x-3)=(\sqrt6x+\sqrt5)(\sqrt6x-\sqrt5)(x-3)\]although it isn't that common to factor across the reals, as opposed to factoring until we get integer or rational coefficients.

OpenStudy (anonymous):

(3x-5)(x+5) this is the final answer I came up with

OpenStudy (anonymous):

I don't think that gives the original expression if you work back.

OpenStudy (anonymous):

6 and 18 are common(by 3) subtract there powers equals 1 so I got 3x-5 then 15 divided by 3 is 5 leaving x-5 .................I may be COMPLETELY wrong, in our weekly work there was one less sets in the work so Im not sure, thanks for the help

OpenStudy (anonymous):

I think if you check my factorization by multiplying, you will find it works back to the original expression.

OpenStudy (anonymous):

k